OpenStudy (anonymous):

\[ \int \frac{x}{x^2+1}dx \]

OpenStudy (phi):

this (almost) looks like u^-1 du

OpenStudy (anonymous):

This smells of parts, but I'm having trouble with \[\int\limits \frac{1}{x^2+1}dx\]

OpenStudy (sirm3d):

that is arctangent

OpenStudy (anonymous):

Although that might not help

OpenStudy (sirm3d):

can you enlighten us with the problem?

OpenStudy (klimenkov):

Do a substitution \(u=x^2, du=2xdx\)

OpenStudy (anonymous):

\[0.5 \ln(u+1)\]

OpenStudy (anonymous):

@sirm3d It's halfway through the DE of my last question

OpenStudy (klimenkov):

Yes. Now make the inverse substitution.

OpenStudy (sirm3d):

post a link @henpen

OpenStudy (sirm3d):

hmm, can we try reduction of order with \[y=v(ix)\]